# Curvature Mapping Method: Mapping Lorentz Force in Orion A

### v2

@@
Magnetic force is a fundamental force in nature. Although widely believed to be important in counterbalancing against collapse in star formation, a clear evaluation of the role of the magnetic field in star formation remains hard to achieve. Past research attempts to evaluate the importance of magnetic forces using diagnostics such as the mass-to-flux ratio, which measures its strength but not how it functions. Since star formation is a complex process and the observed regions have complex structures, mapping the importance of the magnetic field is necessary. We propose a new technique, the Curvature Mapping Method, to evaluate the role of the 2D projected magnetic force by generating maps of the magnetic force derived from polarization observations. This method offers a clear visualization of the contribution of magnetic forces within a region. Its accuracy has been verified in MHD simulations, demonstrating reliable results in most cases. We apply the method to the star formation region of Orion A and provide the first quantitative result where the magnetic force arising from the hourglass magnetic field structure provides support against gravity, which the average Lorentz force is around 310-26 G2 cm-1. The method effectively uses information contained in polarization maps and can be applied to data from surveys to understand the role of the B-field.
@@
The magnetic field plays an important role in star formation (Larson, 1981; McKee & Ostriker, 2007; Li, 2021), whose role in the star formation process is not entirely understood (Crutcher, 2012; Li, 2021; Pattle et al., 2023). Polarization observation (Planck Collaboration et al., 2016; Ward-Thompson et al., 2017; Harper et al., 2018) is the main magnetic field tracer, which can constrain the magnetic field morphology in the plane of the sky (POS). Large radio arrays, such as SMA, and ALMA, provide high-resolution dust polarization data to study magnetic fields of the interstellar medium (ISM) in star formation (Liu et al., 2020, 2024). The next stage is to analyze the information in these data sets to understand physics. The Davis-Chandrasekhar–Fermi (DCF; Davis 1951; Chandrasekhar & Fermi 1953) method is the most commonly used method, where the magnetic field strength is estimated through the dispersion of the polarization angle. The method is based on the assumption that the field lines are straight where the angle dispersion of the field lines is caused by turbulence, and this angle dispersion can be estimated using the dispersion of the polarization angle. However, in many cases, the magnetic field exhibits large-scale coherent structures where the assumption of almost-strain fields as required in the DCF breaks down. The question of how to analyze these data sets becomes an urgent one.
@@
When the field lines are regular, the force is different from assumed in the DCF, where the field line becomes perturbed not to account for the turbulent pressure, but to counterbalance against gravity. In fact, those pinched magnetic field structures are usually interpreted as a balance between magnetic force and gravity in observations(Li et al., 2015; Pattle et al., 2017; Arzoumanian et al., 2021). Following this reasoning, Li et al. 2015 used this force balance to estimate the magnetic field strength by the balance between gravity and the Lorentz force. However, previous studies (Schleuning, 1998; Qiu et al., 2014; Li et al., 2015) only applied this method to some particular cases. We argue that this argument is general and can estimate the Lorentz force in a much larger variety of simulations compared to Li et al. 2015. In this paper, we make the working assumption that the 2D magnetic field, as traced by e.g. the dust polarization observations, reflects the projection of the 3D magnetic field, through which the projected component of the magnetic force on the plane of sky can be estimated through the equation
@@
where the B and are the strength (scalar) and curvature (vector) of magnetic field, the is a modified factor. This approach involves several underlying assumptions: First, the Lorentz force is comparable to the magnetic tension (). Second, the 2D Lorentz force measured in our method reflects the 2D projected component of the 3D Lorentz force. As the discussion of the paper unfolds, one should see that both assumptions hold to some reasonable accuracy. Thus, one can estimate the behavior of the magnetic force using the dust polarization observations, which opens up new approach to the direct analysis of the effect of the magnetic force in the interstellar gas.
@@ 2 Curvature Mapping Method
2.1 Magnetic tension comparable to Lorentz force
@@
where the first term represents magnetic tension and the second term represents magnetic pressure.
The magnetic tension is similar to Lorentz force. Past studies have shown that the energy cascade of MHD turbulence is suppressed by magnetic tension (Grete et al., 2021), suggesting that magnetic tension plays a significant role in the Lorentz force. Comparisons between magnetic tension and the Lorentz force in MHD numerical simulations using the Enzo code (Collins et al., 2012; Burkhart et al., 2015) reveal that the magnetic tension closely resembles the Lorentz force in both magnitude and orientation (details provided in Appendix B). Assuming that magnetic tension approximates the Lorentz force, we can estimate the Lorentz force as:
@@
where represents the magnetic field strength, is the unit vector of the magnetic field, and is the magnetic field curvature vector derived from the magnetic field morphology. Since the magnetic tension, pressure and Lorentz force are similar in the simulation (see Ap. B), the as a weighting factor represents the effect of the magnetic pressure in the Lorentz force. This assumption allows the Lorentz force to be decomposed into two components: a scalar parameter representing magnetic field strength and a vector parameter derived from the magnetic field morphology.
2.2 Adapt observation: from 3D to 2D
@@
The observation of the magnetic field, different from data of the MHD simulation, only has the 2D magnetic field orientation (eg. polarization (Pattle et al., 2017; Chuss et al., 2019; Planck Collaboration et al., 2020a)) in the plane of sky (POS) and the estimation of total magnetic strength (eg. DCF method (Pattle et al., 2023), Zeeman splitting (Crutcher et al., 2010)). Thus, we assume the 2D magnetic tension is close to the Projected component of Lorentz force to measure the projected Lorentz force on POS. According to Eq. 4, this projected Lorentz force depends on the magnetic curvature on POS.
@@
2D curvature vector can be credibly calculated by POS magnetic field orientation in most cases. Curvature is the key parameter of our method, which is calculated by the unit vector of the magnetic field. The POS unit vector can be obtained from magnetic field orientation , the x, y components as the cos and sin. For tracing the projected component of Lorentz force, the magnetic field orientation in POS observed by dust polarization should present the main structure of the 3D magnetic field. When the magnetic field vector along the LOS is less than that of the POS, the POS structure could be the main structure of the whole field, and the offset angle between the LOS vector and the POS plane is less than 45∘. If the magnetic field vector is randomly distributed (Crutcher et al., 2004), the probability of the offset angle between the LOS vector and the POS plane being less than 45∘ is around 71% according to the probability density (). The POS magnetic field structure could present the main magnetic field morphology to calculate the projected curvature and Lorentz force in most cases. The components of the magnetic curvature vector can then be calculated as:
@@
2.3 Tracing the Projected Lorentz force on POS
@@
The projected component of Lorentz force in the plane of the sky (POS) can be measured using the POS magnetic field morphology and magnetic field strength (see Fig. 1). For easy application in observation, the Projected Lorentz force is calculated by the magnetic curvature (2D vector) and total magnetic field strength (scaler) in 2D plane
@@
where is a modified factor including projection effects of the Lorentz force, the effect of magnetic pressure, typically ranging between 1 and 1.5 (detailed in Sect. 3.2).
@@
The POS-projected Lorentz force thus depends on two primary physical parameters: the scalar term (magnetic field strength) and the vector term (magnetic field curvature). The curvature vector can be derived from magnetic field orientations, generally observable through dust polarization (Ward-Thompson et al., 2017; Planck Collaboration et al., 2020b). The total Magnetic field strength, , can be estimated using many methods such as Zeeman splitting (Crutcher et al., 2010), the DCF method (Davis, 1951; Chandrasekhar & Fermi, 1953; Crutcher et al., 2004), the ST method (Skalidis & Tassis, 2021; Skalidis et al., 2021), and so on.
@@
As advances in observations, the projected Lorentz force can be mapped as vectors in the POS, including both its strength and orientation. This technique, known as the *Curvature Mapping Method*, has the added advantage of resolving the Lorentz force orientation over the full range of to . Importantly, curvature is invariant under magnetic field vector direction, ensuring:
@@
Thus, the curvature orientation spans to , unaffected by the polarization-based magnetic field orientation ( to ) observed in the POS. Caution is required in its application under certain conditions, such as when the magnetic field structure is perfectly straight without curved or when magnetic pressure effects dominate significantly, since the assumptions of this method may be less valid in these cases, although these scenarios could be unfrequent (Planck Collaboration et al., 2016; Grete et al., 2021). In most cases, the Curvature Mapping Method effectively measures the projected Lorentz force using the POS magnetic field morphology. The verification of this method is presented in the next section.
3 Verify the Accuracy of the Curvature Mapping Method in Simulation
@@
To verify the accuracy of the curvature mapping method in tracing the POS projected component of Lorentz force, we compare the Lorentz force calculated using the curvature mapping method (referred to as the pseudo Lorentz force) with the projected component of the 3D Lorentz force on the 2D plane (referred to as the projected Lorentz force).
The components of the Lorentz force along the x-, y-, and z-axes are given by:
@@
where the magnetic field vector is . The projected Lorentz force in the x-y plane is:
The pseudo Lorentz force derived using the curvature mapping method is computed from the unit vectors of the magnetic field in the x-y plane and the total magnetic field strength, close to the usage scenario:
where and are the unit vectors defined as and , respectively, in each x-y plane, and is the total magnetic field strength. The modified factor is temporarily assumed to be 1. The x and y components of the pseudo Lorentz force are:
In this way, we simulate random observational perspectives using MHD simulations with 2563 data points to verify the accuracy of the Curvature Mapping Method in determining both the orientation and strength of the Lorentz force.
3.1 Comparison of Orientations

### v3

@@
Magnetic forces are critical in regulating the star formation process, yet spatially resolving their dynamical role remains challenging. Traditional diagnostics like the mass-to-flux ratio measure field strength but lack directional information on how the force functions. We propose the Curvature Mapping Method (CMM) to reconstruct the 2D projected Lorentz force () vector field from polarization observations. The method assumes that the magnetic field is deformed by external forces such as gravity, marking a transition from a force-free to a forced regime. We validate the method using MHD simulations, confirming its reliability specifically in dense molecular cloud cores where the magnetic field structure is well-resolved and magnetic forces are dynamically significant. We apply CMM to the OMC-1 region and compare the derived map with a spatially resolved gravitational force map () calculated via the Poisson equation. We identify a clear bimodal behavior: in the diffuse envelope ( cm-2), the magnetic force is statistically uncorrelated with gravity; however, along the dense filamentary ridge, vectors become systematically anti-parallel to gravity. The magnetic support ratio, , approaches unity in the filament backbone, providing direct evidence that magnetic fields provide substantial support against gravitational collapse, thereby regulating the fragmentation and star formation rate. The method effectively uses information contained in polarization maps and can be applied to data from surveys to understand the role of the B-field.
@@
The magnetic field plays an important role in the interstellar medium (ISM) and molecular clouds (Larson, 1981; McKee and Ostriker, 2007; Li, 2021), whose role in the star formation process is not entirely understood (Crutcher, 2012; Li, 2021; Pattle et al., 2023). Thermal dust polarization, arising from aspherical dust grains aligned with magnetic field lines (Planck Collaboration et al., 2016; Ward-Thompson et al., 2017; Harper et al., 2018), is the main magnetic field tracer, which can constrain the magnetic field morphology in the plane of the sky (POS). Large radio arrays, such as SMA and ALMA, provide high-resolution dust polarization data to study magnetic fields of the interstellar medium (ISM) in star formation (Liu et al., 2020, 2024). The next stage is to analyze the information in these data sets to quantify the dynamical properties of magnetic fields, such as their field strength and their role in counteracting self-gravity. Various techniques have been proposed to measure magnetic fields in the ISM, such as the Zeeman effect and Faraday rotation (Crutcher, 2012; 2025ApJ...978L..31R). However, for dust polarization observations in molecular clouds, the Davis-Chandrasekhar–Fermi (DCF; Davis 1951; Chandrasekhar and Fermi 1953) method is the most commonly used method, where the magnetic field strength is estimated through the dispersion of the polarization angle. The method is based on the assumption that the field lines are straight where the angle dispersion of the field lines is caused by turbulence, and this angle dispersion can be estimated using the dispersion of the polarization angle. However, in many cases, the magnetic field exhibits large-scale coherent structures (2006Sci...313..812G; Qiu et al., 2014; Li et al., 2015) where the assumption of almost-straight fields as required in the DCF breaks down. The question of how to analyze these data sets becomes an urgent one.
@@
In contrast to the DCF method, which attributes magnetic field distortions to turbulent motions, the curvature of ordered field lines serves a different dynamical role: generating magnetic tension to counterbalance gravity. In fact, those pinched magnetic field structures are usually interpreted as a balance between magnetic force and gravity in observations (Li et al., 2015; Pattle et al., 2017; Arzoumanian et al., 2021). Following this reasoning, Li et al. (2015) used this force balance to estimate the magnetic field strength by the balance between gravity and the Lorentz force. However, previous studies (Schleuning, 1998; 2012ApJ...747...79K; Qiu et al., 2014; Li et al., 2015) only applied this method to some particular cases. We argue that this argument is general and can estimate the Lorentz force in a much larger variety of situations compared to Li et al. 2015. In this paper, we make the working assumption that the observed curvature arises from the resistance of the magnetic field to external compression (e.g., self-gravity), moving the system away from a force-free state () toward a dynamic equilibrium. The 2D magnetic field, as traced by e.g. the dust polarization observations, reflects the projection of the 3D magnetic field, through which the projected component of the magnetic force on the plane of sky can be estimated through the equation:
@@
where and are the strength scalar and curvature vector of magnetic field, and is a numerical factor. This approach involves two underlying assumptions: First, the Lorentz force is comparable to the magnetic tension (). Second, the 2D Lorentz force measured in our method reflects the 2D projected component of the 3D Lorentz force. As the discussion of the paper unfolds, one should see that both assumptions hold to some reasonable accuracy. Thus, one can estimate the behavior of the magnetic force using the dust polarization observations, which opens up new approach to the direct analysis of the effect of the magnetic force in the interstellar gas.
@@ 2 Curvature Mapping Method
2.1 Magnetic Tension as a proxy for the Lorentz Force
@@
where the first term represents magnetic tension and the second term represents magnetic pressure force.
Magnetic tension, acting as the restoring force against field line bending, is a fundamental component of the Lorentz force. Direct comparisons between magnetic tension and the Lorentz force in MHD numerical simulations using the Enzo code (Collins et al., 2012; Burkhart et al., 2015) reveal that the magnetic tension closely resembles the Lorentz force in both magnitude and orientation (details provided in Section 3.2). While magnetic pressure force contributes significantly to the force magnitude (comparable to tension), MHD simulations (see Sect. 3.2) reveal that the magnetic tension vector aligns much more closely with the total Lorentz force orientation in dense gas than the pressure gradient does. This factor accounts for the combined effects of the magnetic pressure force and the 3D-to-2D projection geometry, bridging the gap between the projected tension and the total 3D Lorentz force. The Lorentz force is estimated as:
@@
where represents the magnetic field strength, is the unit vector of the magnetic field, and is the magnetic field curvature vector derived from the magnetic field morphology. Since the magnetic tension, pressure force and Lorentz force are similar in the simulation (see Sect. 3.2), the weighting factor is introduced to account for the contribution of the magnetic pressure force (which is not directly traced by curvature) to the total Lorentz force magnitude. The Tension Ansatz: using tension to represent the Lorentz force works in the kinematic regime, but not the force-free region, and this corresponds to the high-density part of the ISM (Li and Zhao, 2025) where the magnetic force is forced rather than force-free. This assumption allows the Lorentz force to be decomposed into two components: a scalar parameter representing magnetic field strength and a vector parameter derived from the magnetic field morphology.
2.2 Observational Adaptation: From 3D to 2D
@@
The observation of the magnetic field, different from data of the MHD simulation, only has the 2D magnetic field orientation (Pattle et al., 2017; Chuss et al., 2019; Planck Collaboration et al., 2020a, e.g., polarization) in the plane of sky (POS) and the estimation of total magnetic strength (Pattle et al., 2023; Crutcher et al., 2010, e.g., DCF method, Zeeman splitting). Thus, we assume the 2D magnetic tension is close to the Projected component of Lorentz force to measure the projected Lorentz force on POS. The robustness of this assumption is quantitatively verified in Section 3.3, where we demonstrate that the derived 2D force orientation closely tracks the true projected Lorentz force in dense environments. According to Eq. 4, this projected Lorentz force depends on the magnetic curvature on POS.
@@
The 2D curvature vector can be credibly calculated by POS magnetic field orientation in most cases. Curvature is the key parameter of our method, which is calculated by the unit vector of the magnetic field. The POS unit vector can be obtained from magnetic field orientation , the x, y components as the cos and sin. To trace the projected component of the Lorentz force, the magnetic field orientation observed in the POS via dust polarization is assumed to represent the main structure of the 3D magnetic field. Assuming a random distribution of magnetic field orientations (Crutcher et al., 2004), the probability of the offset angle () between the 3D field vector and the POS being less than 45∘ is approximately 71 (based on the probability density ). This implies that in most cases, the POS structure captures the dominant magnetic field morphology. However, projection effects inherently introduce uncertainty into individual magnitude estimates, which scale with . To address this, we quantitatively evaluated the impact of projection in Section 3.3. As shown in Fig. 5, we analyzed the ratio of the derived force to the true 3D Lorentz force. The results demonstrate that in dense gas, this ratio distribution is well-constrained. This confirms that the statistical correction factor, , effectively accounts for the projection geometry.The POS magnetic field structure could present the main magnetic field morphology to calculate the projected curvature and Lorentz force in most cases. The components of the magnetic curvature vector can then be calculated as:
@@
2.3 Tracing the Projected Lorentz Force
@@
The projected component of Lorentz force in the plane of the sky (POS) can be measured using the POS magnetic field morphology and magnetic field strength (see Fig. 1). For easy application in observation, the Projected Lorentz force is calculated by the magnetic curvature and total magnetic field strength in the 2D plane
@@
where is a modified factor including projection effects of the Lorentz force, the effect of magnetic pressure force, typically ranging between 1 and 1.5 (detailed in Sect. 3.3.2).
@@
The POS-projected Lorentz force thus depends on two primary physical parameters: the scalar term (magnetic field strength) and the vector term (magnetic field curvature). The curvature vector can be derived from magnetic field orientations, generally observable through dust polarization (Ward-Thompson et al., 2017; Planck Collaboration et al., 2020b). The total Magnetic field strength, , can be estimated using many methods such as Zeeman splitting (Crutcher et al., 2010), the DCF method (Davis, 1951; Chandrasekhar and Fermi, 1953; Crutcher et al., 2004), the ST method (Skalidis and Tassis, 2021; Skalidis et al., 2021), and so on.
@@
As advances in observations, the projected Lorentz force can be mapped as vectors in the POS, including both its strength and orientation. This technique, the *Curvature Mapping Method*, has the added advantage of resolving the Lorentz force orientation over the full range of to . Importantly, curvature is invariant under magnetic field vector direction, ensuring:
@@
Thus, the curvature orientation spans to , unaffected by the polarization-based magnetic field orientation ( to ) observed in the POS. Caution is required in its application under certain conditions, such as when the magnetic field structure is perfectly straight without curved or when magnetic pressure force effects dominate significantly, since the assumptions of this method may be less valid in these cases, although these scenarios could be infrequent (Planck Collaboration et al., 2016; Grete et al., 2021). In most cases, the Curvature Mapping Method effectively measures the projected Lorentz force using the POS magnetic field morphology. The verification of this method is presented in the next section.
3 Verification via MHD Simulations
3.1 Simulations
The numerical simulation of molecular cloud selected in this work is identified within the three simulations applied by the constrained transport MHD option in Enzo (MHDCT) code (Collins et al., 2010; Burkhart et al., 2015). The simulation conducted in this study analyzed the impact of self-gravity and magnetic fields on supersonic turbulence in isothermal molecular clouds, using high-resolution simulations and adaptive mesh refinement techniques. which detail in shown in (Collins et al., 2012; Burkhart et al., 2015, 2020). The Enzo simulation with three dimensionless parameters provides us with massive information on physical processes:
where is the rms velocity fluctuation, cs is sonic speed, is mean density, and is the mean magnetic field strength. With the same sonic Mach number and viral parameter in initial conditions, the difference only exists in initial magnetic pressure force with various B-field strength, The as 0.2, 2, and 20 present the strong B-field state, medium B-field, and weak B-field state. The clouds are in three different evolutionary stages started with self-gravity existing at 0.6 ( Myr), where represents 1.1 Myr (Collins et al., 2012; Burkhart et al., 2015). Due to the short timescale of evolution, the gravitational collapse could barely affect the B- relation in the simulation The size of these molecular clouds is around 4.63 pc (2563 pixels), with one-pixel size of approximately 0.018 pc (Burkhart et al., 2015).
In this Enzo simulation, the thermal energy is located at the low state ( = 9). Their timescale of evolution stage is 0.6 , around 0.76 Myr, which the timescale starts from when interstellar media collapse. Due to the short evolutionary time ( 3.85 10-21 g cm-3, or 821 cm-3), the gravitational energy less affects the system in simulation. With the similar mean density in these simulations, the different (0.2, 2, 20) present the different magnetic field states: strong B-field, stronger B-field (weaker than strong B-field but stronger than weak B-field), and weak B-field. The Alfvén Mach number in various simulations can be calculated by the energy ratio between the total magnetic energy and kinetic energy (Zhao et al., 2024):
3.2 Validity of the Physical Assumption: Magnetic Tension vs. Pressure Force
We employ the simulation ( Myr) to validate the physical assumptions of the Curvature Mapping Method, specifically comparing the roles of magnetic tension () and magnetic pressure force () within the total Lorentz force (). We first evaluate the global vector similarity in 3D space:
where the is the vector of Lorentz force and could be magnetic tension or pressure force. Using the similarity metric , we find that both tension and pressure force statistically resemble the Lorentz force structure (see Fig. 3). After correcting for projection effects using the probability density function (see Sect. B), the mean projected offset angle between magnetic tension and the total Lorentz force is approximately , while the offset for magnetic pressure force is around (Fig. 3). This indicates that, on a global scale, for initial condition similar to that of the molecular cloud, the magnetic pressure force and tension align such that the direction of either, e.g. the magnetic tension, can be used to estimate the direction of the Lorentz force
To rigorously assess the physical validity of our method, we investigated the density dependence of these correlations using the full 3D vector fields. Figure 2 presents the statistical comparison of magnitude ratios and angular offsets as a function of gas density. Regarding magnitude, both magnetic tension and pressure force remain comparable to the total Lorentz force across the entire density range. This confirms that the magnetic pressure force contribution is non-negligible in strength, justifying the introduction of the weighting factor (Eq.,8) to recover the total force magnitude. The distribution of angular offsets are both below 30∘ in the dense gas regime (). Note that produce the distribution of angular offset, we have taken corrected for the geometric projection effect (see Ap. B) to reveal the true physical alignment. The magnetic tension vector maintains a tight alignment with the total Lorentz force, characterized by a corrected mean offset of (Fig. 2c). In contrast, the magnetic pressure force exhibits significantly larger misalignment () and broader dispersion in the same regime (Fig. 2d).
The choice of curvature (the geometric indicator of magnetic tension) as a tracer is driven not only by this alignment advantage but also by observational feasibility. The magnetic pressure force depends on the spatial gradient of the field strength (), which is notoriously difficult to constrain in observations (Davis, 1951; Chandrasekhar and Fermi, 1953; Skalidis and Tassis, 2021; Liu et al., 2022). In contrast, magnetic tension is geometrically determined by the field curvature (), a vector quantity that can be robustly derived from the magnetic field morphology traced by dust polarization (see Sect. 2). Therefore, combining the intrinsic directional alignment with observational accessibility, magnetic curvature stands out as the optimal proxy for mapping the Lorentz force orientation in molecular clouds. We also verified the analysis by projecting the simulation data along the x-z and y-z planes. The statistical results remain consistent across these orthogonal projections, confirming the robustness of the method.
3.3 Accuracy of Curvature Mapping Method in Determining Lorentz Force
@@
To verify the accuracy of the Curvature Mapping Method in tracing the POS projected component of the Lorentz force, we compare the force derived from our method (referred to as Sect. 2) with the true projected component of the 3D Lorentz force calculated directly from the simulation.
The components of the true 3D Lorentz force are given by:
The projected Lorentz force in the - plane is its projection:
The CMM-derived projected Lorentz force is computed following the observational procedure, utilizing only the unit vector of the magnetic field in the - plane and the total magnetic field strength:
@@
where is the unit vector of the magnetic field in the - plane (), and (Crutcher et al., 2004) ( is the projected field strength). The and components of the CMM-derived projected Lorentz force are calculated as:
We analyzed 2D slices along the -axis from the MHD simulations to simulate random observational perspectives and evaluate the method’s performance. We also verified the analysis by projecting the simulation data along the x-z and y-z planes. The statistical results remain consistent across these orthogonal projections, confirming the robustness of the method.
3.3.1 Comparison of Directions